OpenStudy (anonymous):

What is the value of 19 + 5a^4 – 3b^2 when a = 2 and b = −4?

jagr2713 (jagr2713):

Hey what do you think we do first?

OpenStudy (anonymous):

hi @jagr2713 :D long time no see

OpenStudy (anonymous):

ok we multiply the exponets

OpenStudy (anonymous):

right?

jagr2713 (jagr2713):

correct spo what do we do with the 2^4

OpenStudy (anonymous):

Multiply the exponents? Do you mean apply the exponents?

OpenStudy (anonymous):

i think????

OpenStudy (anonymous):

do you guys have khans academy account?

OpenStudy (anonymous):

Yeah, I do. But how is that relevant to the question?

OpenStudy (anonymous):

so you can tell me where do i get videos from math n stuff

jagr2713 (jagr2713):

So we get 16 so we have 19+5*16-3(-4)^2 since the powers of 2 is even, the result will be positive 19+5*16-3*4^2 Now we sumplify 4^2 to 16 19+5*16-3*16 Can you use PEMDAS fro here

OpenStudy (anonymous):

256???

OpenStudy (anonymous):

...

jagr2713 (jagr2713):

no we simplify 5*16 to get 80 19+80-3*16 Now 3*16=48 19+80-48 solve from here

OpenStudy (anonymous):

\[\large 19~+~(5 \times 16)~-~(3 \times 16)\]

OpenStudy (anonymous):

51 (reacts to kid who says 21)

jagr2713 (jagr2713):

Use PEMDAS

jagr2713 (jagr2713):

Correct is 51!

OpenStudy (anonymous):

thx :D

jagr2713 (jagr2713):

Glad to had helped today!